NEAR INFRARED LASER SPECTROSCOPY OF SCANDIUM MONOBROMIDE

\underline{YE XIA}, A. S.-C. CHEUNG, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong.; ZHENWU LIAO, MEI YANG, MAN-CHOR CHAN, Department of Chemistry, The Chinese University of Hong Kong, Shatin, N.T., Hong Kong.

High resolution laser spectrum of scandium monobromide (ScBr) between 787 and 845 nm has been investigated using the technique of laser vaporization/reaction with free jet expansion and laser induced fluorescence spectroscopy. ScBr was produced by reacting laser vaporized Sc atoms with ethyl bromide (C_{2}H_{5}Br). Spectra of six vibrational bands of both Sc^{79}Br and Sc^{81}Br isotopomers of the C^{1}Σ^{+} - X^{1}Σ^{+} transition and seven vibrational bands of the e^{3}Δ - a^{3}Δ transition were obtained and analyzed. Least-squares fit of the measured line positions for the singlet transitions yielded accurate molecular constants for the v = 0 3 levels of the C^{1}Σ^{+} state and the v = 0 2 levels of the X^{1}Σ^{+} state. Similar leastsquares fit for the triplet transitions yielded molecular constants for the v = 0 2 levels of both e^{3}Δ and a^{3}Δ states. The equilibrium bond length, r_{0}, of the a^{3}Δ state has been determined to be 2.4789 Å. Financial support from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. HKU 701008P) is gratefully acknowledged